OpenStudy (anonymous):

If f(x) = 2x + 2, find f(5).

OpenStudy (***[isuru]***):

hi, This is really an easy one u just have to substitute "5" instead of x in F(x) F(x) = 2x +2 = (2*5) + 2 = 12

OpenStudy (anonymous):

Ok I Got That Right Then